# QDM: Quadtree-Based Region-Adaptive Sparse Diffusion Models for Efficient Image Super-Resolution

### v1

@@ Appendix C Additional Experimental Results
Table 6: Quantitative comparisons of the proposed QDM-L equipped with different threshold, on the benchmark LSDIR-Test.
@@
Results on RealSet80 [44].
Since the RealSet80 benchmark lacks reference ground-truth images, we evaluate model performance exclusively using non-reference metrics. As shown in Table 7, QDM-L achieves comparable performance to other training-from-scratch methods.
Table 7: Quantitative comparisons of various methods on RealSet80. The best and second-best results are bold and underlined, respectively.
Methods Metrics
[…]
Pretrained-Prior-Based Methods
StableSR-50 [33] 0.6214 62.7613
DiffBIR-50 [19] 0.7404 67.9806
SeeSR-50 [38] 0.7114 69.7658
OSEDiff-1 [39] 0.7093 68.8022
InvSR-1 [42] 0.7291 69.8055
SinSR-1 [36] 0.7228 64.0573
Training-from-Scratch Methods
BSRGAN [45] 0.6263 66.6288
RealESRGAN [35] 0.6189 64.4957
SwinIR [17] 0.6050 64.6984
LDM-50 [25] 0.5511 55.8246
ResShift-15 [44] 0.6564 62.0441
ResShift-4 [43] 0.6515 61.7967
QDM-L-15 0.6643 61.7969
Ablation study of quadtree threshold on LSDIR-Test.
We also conducted an ablation study on the quadtree threshold using real-world benchmarks. However, real-world images often contain significant noise and lack large homogeneous regions, making them less suited for adaptive computation. As shown in Table 3, at s=0 (lossless mode), QDM-L achieves performance comparable to the full-mask baseline but exhibits minimal computation reduction, as the mask remains nearly dense. Increasing s to $0.09$ reduces computation by $3.22\%$ compared to the baseline, but this comes at a significant performance cost. This suggests that higher quadtree thresholds are not well-suited for super-resolution tasks involving noisy and complex low-resolution inputs.
Qualitative comparisons of different methods.
We also include more qualitative comparison of different methods in Figure 7 and Figure 8.

### v2

@@ Appendix C Additional Experimental Results
Table 7: Downstream-only and prior-based methods on Med-8
Metric Downstream-only Prior-based QDM-L
DiffBIR SeeSR
[…]
Table 8: Quantitative comparisons of the proposed QDM-L equipped with different thresholds, on the benchmark LSDIR-Test.
@@
Downstream-only Design on Med-SR8. We designed the dual‐stream architecture to leverage the complementary strengths of both branches. The downstream branch focuses on local, detail-rich regions while the upstream branch captures global context and applies minimum refinement on homogeneous regions. Here, we conducted experiments to validate that without the upstream branch, the model cannot capture the global view and is prone to overlook large‐scale structure, leading to degraded performance, as shown in Table 7. More importantly, a downstream‐only design lacks the ability to allocate computation adaptively across different regions of the image.
Pretrained-Prior-Based Methods on Med-SR8. We’ve included detailed results on natural images comparing to several prior-based methods in Table 1. To test their generalization abilities, we evaluated two methods (DiffBIR and SeeSR) on the Med-8 benchmark. Neither method yielded satisfactory reconstructions in the medical domain, with notably lower performance compared to QDM (Table 7). These results highlight a substantial generalization gap, suggesting that such methods may require significant domain-specific data and fine-tuning.
Ablation study of quadtree threshold on LSDIR-Test. We also conducted an ablation study on the quadtree threshold using real-world benchmarks. However, real-world images often contain significant noise and lack large homogeneous regions, making them less suited for adaptive computation. As shown in Table 4, at s=0 (lossless mode), QDM-L achieves performance identical to the full-mask baseline but exhibits minimal computation reduction, as the mask remains nearly dense. Increasing $s$ to $0.09$ reduces computation by $3.22\%$ compared to the baseline, but this comes at a significant performance cost. This suggests that higher quadtree thresholds are not well-suited for super-resolution tasks involving noisy and complex low-resolution inputs.
Appendix D Qualitative comparisons of different methods.
We also include more qualitative comparisons of different methods in Figure 7, Figure 8 and Figure 9.
@@
Figure 9: Qualitative comparison of CECTs in the tumor region.